Shift report 20030315 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Mar 15 2003 - 08:21:22 EST

  • Next message: Apache: "Shift report 20030315 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Mar 15 08:21:22 2003, Shift: 00:00-08:00
    Supervisor: CEJ, with: 
    Log book#14 pages 62 - 62
    
    Shift Summary:
    ------------------------------------------------------------
    1) Beam untill 01:00. Good store from 03:00, lost due to 
       yellow QLI at 05:20.
    
    2) Ped, TPC ped and DVM runs 8385, 8386 and 8387
       MRS@90(B1050), FS@12(A1/4):
       8383 Trig2 3K, trig3 33K,
       8384 Trig2 1.6K trig3 17K (T3 & T4 problems)
       MRS@90(B1050), FS@12(A1/2):
       8388 Trig2 1.7K, trig3 153K (T3 problem)
       8388 Trig2 0.18K, trig3 16K (T3 problem)
    
    3) T3 and T4 field and cathode tripped. T4 was brought 
       back, but not T3. After turning on field and cathodes  
       again, the anodes continued tripping (see E-log for  
       details). DVM3 gave a funny shaped spectrum, but T1 
       looked ok.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Mar 15 2003 - 08:25:13 EST